# Textoon: Generating Vivid 2D Cartoon Characters from Text Descriptions

### v1

@@ 1 Introduction
Cartoon characters are fictional beings often often distinguished by their adorable appearances and vivid colors. They are widely used in films, games, social media, and advertising. Cartoon characters can be crafted in both 2D and 3D styles. While 3D animations provide greater freedom for creation and control, they typically come with higher costs and depend on rendering engines. Conversely, creating 2D cartoon digital characters is more straightforward and efficient, making them especially suitable for devices with limited processing power, such as mobile phones and web applications. In the realm of 2D cartoon characters, Live2D[3] has emerged as a leading standard for delivering real-time interactive performance.

### v2

@@ 1 Introduction
Cartoon characters are fictional beings often distinguished by their adorable appearances and vivid colors. They are widely used in films, games, social media, and advertising. Cartoon characters can be crafted in both 2D and 3D styles. While 3D animations provide greater freedom for creation and control, they typically come with higher costs and depend on rendering engines. Conversely, creating 2D cartoon digital characters is more straightforward and efficient, making them especially suitable for devices with limited processing power, such as mobile phones and web applications. In the realm of 2D cartoon characters, Live2D[5] has emerged as a leading standard for delivering real-time interactive performance.
@@ 3.7 Animation
For the speech-driven facial animation, we designed a network that takes speech as input and outputs blendshape coefficients to achieve vivid facial expressions. As shown in Fig. 10, the network follows an encoder-decoder architecture. We adopt the state-of-the-art pre-trained speech model Wav2Vec-XLSR [2, 3] for the audio encoder. The features extracted from raw audio waveform are combined with style features and fed into a transformer-based decoder, which outputs stylized blendshape coefficients. The Fig. 11 also illustrates the facial animation generated from English and Chinese speeches.
@@
Figure 10: Animation framework. The Wav2Vec-XLSR [2, 3] is used to extract audio features from a speech signal. The audio features, combined with style features, are fed into a transformer to output stylized blendshape coefficients.
Figure 11: Visualization for audio-driven Live-2D face animation.